solving the equation with radicals
w=square root 2w+15

Answer by checkley77(12844)

w=sqrt(2w+15) square both sides.
w^2=2w+15
w^2-2w-15=0
(w-5)(w+3)=0
w-5=0
w=5 ans.
w+3=0
w=-3 ans.